# Invariant subspaces and PCA in nearly matrix multiplication time

### v3

@@ B.1 Spectral gaps with diagonalization
The straightforward approach to compute the desired spectral gaps is to iteratively compute $\boldsymbol{\mathrm{S}}^{-1}\boldsymbol{\mathrm{H}}$ and diagonalize it, until the eigenvalues are well approximated. If we compute $\boldsymbol{\mathrm{A}}\leftarrow\mathsf{MM}(\mathsf{INV}(\boldsymbol{\mathrm{S}}),\boldsymbol{\mathrm{H}})$, then we can write $\boldsymbol{\mathrm{A}}=\boldsymbol{\mathrm{S}}^{-1}\boldsymbol{\mathrm{H}}+\boldsymbol{\mathrm{E}}$, for some error matrix $\boldsymbol{\mathrm{E}}$. The next step is to approximate $\operatorname{gap}_{k}(\boldsymbol{\mathrm{S}}^{-1}\boldsymbol{\mathrm{H}})$ and $\mu_{k}(\boldsymbol{\mathrm{S}}^{-1}\boldsymbol{\mathrm{H}})$. For this one could use the recent state-of-the-art backward-approximate diagonalization algorithm of [BGVKS22]. We highlight that a similar result can be obtained by using the more recent pencil diagonalization algorithm of [DDS23]. The authors the latter mention that their algorithm should be “more numerically stable” than $\mathsf{EIG}$ of [BGVKS22], since it uses an inverse-free iteration internally, and they do provide strong theoretical and experimental evidence for this statement. However, as there is no formal, end-to-end proof for the bit complexity at the time of this writing, we choose to compare against $\mathsf{EIG}$ of [BGVKS22]. Note that, any improvements on the bit requirements for the matrix sign function, directly provide the same improvements for our main Theorems 1.1 and 3.1.
Theorem B.1 ($\mathsf{EIG}$, imported Theorem 1.6 from [BGVKS22]).
@@
Applying $\mathsf{EIG}$ on $\boldsymbol{\mathrm{A}}$ we can obtain a backward-approximate diagonalization. But we are not finished yet, since we are interested in each individual eigenvalue. To translate the backward error to a forward error for the eigenvalues, we state the following Corollary, which is an immediate consequence of Kahan’s inequality (Fact A.1).
@@ Appendix C Analysis of Cholesky
Algorithm 2 $\mathsf{CHOLESKY}(\boldsymbol{\mathrm{M}})$.
1:
$\mathsf{CHOLESKY}$
3:$\boldsymbol{\mathrm{M}}$ is Hermitian and positive-definite, $\textbf{{u}}\leq\epsilon\frac{1}{c_{1}n^{c_{2}}\kappa(\boldsymbol{\mathrm{M}})^{c_{3}\log(n)}}$ for some constants $c_{1},c_{2},c_{3}$ and $\epsilon\in(0,1)$.
5:if $n=1$ :
6:    return $\boldsymbol{\mathrm{L}}=\sqrt{\boldsymbol{\mathrm{M}}_{1,1}}$.
7:else
9:    $\boldsymbol{\mathrm{B}}\boldsymbol{\mathrm{A}}i\leftarrow\mathsf{MM}\left(\boldsymbol{\mathrm{B}},\mathsf{INV}\left(\boldsymbol{\mathrm{A}}\right).\right)$triangleright $\boldsymbol{\mathrm{B}}\boldsymbol{\mathrm{A}}i=\boldsymbol{\mathrm{B}}\left(\boldsymbol{\mathrm{A}}^{-1}+\boldsymbol{\mathrm{E}}_{1}^{\mathsf{INV}}\right)+\boldsymbol{\mathrm{E}}_{2}^{\mathsf{MM}}=\boldsymbol{\mathrm{B}}\boldsymbol{\mathrm{A}}^{-1}+\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{B}}\boldsymbol{\mathrm{A}}i}.$
@@ C.1 Error analysis
which means that it suffices to bound the individual terms in the sum. The first term is the error of the first recursive call. Following the notation of [DDH07], let us denote $\boldsymbol{\mathsf{err}}(n)$ the norm-wise error of the algorithm for size $n$. We can then write
@@ C.2 Maintaining positive-definiteness, norms, and condition numbers throughout the recursion
C.3 Final backward-approximation bounds and proof of Theorem
Having safeguarded the possibility of a breakdown due to loss of positive-definiteness, we can now revisit the bounds of Lemma C.1.
@@ C.3 Final backward-approximation bounds and proof of Theorem
For the proof we assume the bound for $\textbf{{u}}\leq\textbf{{u}}_{++}$ from Inequality (13). Combining the inequalities for a single recursion step from Lemma C.2 with the results of the previous section and the bound for u, we have the following inequalities (recall once more that $\boldsymbol{\mathrm{M}}$ is the input of the algorithm for the specific recursive call, while $\boldsymbol{\mathrm{M}}_{0}$ is the matrix in the original call):
@@ C.4 Reducing the GEP to a regular Hermitian eigenproblem
Given $\boldsymbol{\mathrm{H}}\in\mathbb{H}^{n}$, $\boldsymbol{\mathrm{S}}\in\mathbb{H}^{n}_{++}$, and $\epsilon\in(0,1),$ Algorithm 3 executes $O(T_{\mathsf{MM}}(n)$ floating point operations and returns a matrix $\widetilde{\boldsymbol{\mathrm{H}}}\leftarrow\mathsf{REDUCE}(\boldsymbol{\mathrm{H}},\boldsymbol{\mathrm{S}},\epsilon)$, such that, if the machine precision satisfies
@@ C.5 Counterexample for floating point LU
In this section provide a counter-example which illustrates why the backward-stable LU factorization algorithm of [DDH07] is not sufficient to obtain a Cholesky factor for a Hermitian postive-definite matrix. We apply one-by-one the steps of the $\mathsf{LUR}$ algorithm of [DDH07] on the matrix:
@@ Appendix D Regularization with diagonal disorder
We now arrive to the more interesting part of the analysis. In Section 2 we argued that if we know the gap and the midpoint then the sign function can yield the desired spectral projector. To compute the gap and midpoint efficiently, we will take advantage of the symmetry induced by the Cholesky-based reduction, and use the Wegner estimate [Weg81] for Hermitian diagonal perturbations to regularize the problem in the spirit of smoothed analysis. We use the following variant of the Wegner estimate.
@@
This directly comes from the well-known result of Wegner [Weg81], which states that that for any interval $I\subset\mathbb{R}$
@@
Similar results can be obtained for other classes of random matrices $\boldsymbol{\mathrm{G}}$. Such an example is the Gaussian Unitary Ensemble (GUE) [APS+17], in which case $\boldsymbol{\mathrm{G}}$ is dense, but it is invariant under rotations which might be useful for other applications.
@@ E.2 Approximate midpoint and gap
Theorem E.1 ($\mathsf{GAP}$).
@@
arithmetic operations using $O\left(\log(n)\left(\log(\tfrac{1}{\delta\epsilon\operatorname{gap}_{k}})\log^{3}(\tfrac{n}{\delta\epsilon\operatorname{gap}_{k}})+\log(\kappa(\boldsymbol{\mathrm{S}}))\right)\right)$ bits, where $\lambda_{i}$ are the eigenvalues of $(\boldsymbol{\mathrm{H}},\boldsymbol{\mathrm{S}})$. If $\kappa(\boldsymbol{\mathrm{S}})$ is unknown, additional $O(T_{\mathsf{MM}}(n)\log(\tfrac{n\kappa(\boldsymbol{\mathrm{S}})}{\delta})\log(\kappa(\boldsymbol{\mathrm{S}})))$ floating point operations and $O(\log(n)\log(\tfrac{\kappa(\boldsymbol{\mathrm{S}})}{\delta})\log^{3}(\tfrac{n\kappa(\boldsymbol{\mathrm{S}})}{\delta}))$ bits are required to compute it.
@@
$O\left(T_{\mathsf{MM}}(n)\log(\tfrac{n}{\gamma\delta})\right)$
arithmetic operations using $O\left(\log(n)\log(\tfrac{1}{\gamma\delta})\log^{3}(\tfrac{n}{\gamma\delta})\right)$ bits. The last iteration where $\gamma_{m}=\Theta(\epsilon\operatorname{gap}_{k})$ and $\delta_{m}=\Theta(\delta_{0}\epsilon\operatorname{gap}_{k})$ is the most expensive, which gives a total of
$O\left(T_{\mathsf{MM}}(n)\log(\tfrac{n}{\delta_{0}\epsilon\operatorname{gap}_{k}})\log(\tfrac{1}{\epsilon\operatorname{gap}_{k}})\right)$
@@
$O\left(\log(n)\log(\tfrac{1}{\delta_{0}\epsilon\operatorname{gap}_{k}})\log^{3}(\tfrac{n}{\delta_{0}\epsilon\operatorname{gap}_{k}})\right)$
@@
To be able to call $\mathsf{REDUCE}$ in every step, we need to approximate $\kappa(\boldsymbol{\mathrm{S}})$ in order to set the machine precision appropriately. For this we can use Corollary E.1, which is detailed in the next section, and it returns $\widetilde{\kappa}\in\Theta(\kappa(\boldsymbol{\mathrm{S}}))$ with probability $1-\delta$. It requires $O(T_{\mathsf{MM}}(n)\log(\tfrac{n\kappa(\boldsymbol{\mathrm{S}})}{\delta})\log(\kappa(\boldsymbol{\mathrm{S}})))$ floating point operations using $O(\log(n)\log(\tfrac{\kappa(\boldsymbol{\mathrm{S}})}{\delta})\log^{3}(\tfrac{n\kappa(\boldsymbol{\mathrm{S}})}{\delta}))$ bits. ∎
@@ E.3 Singular values, singular gaps, and condition number
$O\left(T_{\mathsf{MM}}(n)\left(\tfrac{n}{m}+\log(\tfrac{n}{\delta\epsilon\sigma_{k}})\right)\log(\tfrac{1}{\epsilon\sigma_{k}})\right)$
floating point operations and requires $O\left(\log(n)\log(\tfrac{1}{\delta\epsilon\sigma_{k}})\log^{3}(\tfrac{n}{\delta\epsilon\sigma_{k}})\right)$ bits.
@@
We will solve the problem with counting queries, similar to Theorem 3.1, halving $\gamma$ and $\delta$ in every iteration. We start with $\gamma_{0}=1/8$ and $\delta_{0}=\delta/2$. First, we divide $\boldsymbol{\mathrm{A}}$ by two, to ensure that $\|\boldsymbol{\mathrm{A}}\|\leq 1/2$. Then we construct $\widetilde{\boldsymbol{\mathrm{A}}}=\mathsf{HERM}(\mathsf{MM}(\boldsymbol{\mathrm{A}},\boldsymbol{\mathrm{A}}^{*}))=\boldsymbol{\mathrm{A}}\boldsymbol{\mathrm{A}}^{*}/4+\boldsymbol{\mathrm{E}}^{\mathsf{MM}}$. To perform this rectangular multiplication in practice, we partion $\boldsymbol{\mathrm{A}}$ in $d/n$ blocks of size $n\times n$ each, perform the individual multiplications and sum the results. We can ensure that $\|\boldsymbol{\mathrm{E}}^{\mathsf{MM}}\|\leq\gamma_{0}/4$ if we use $O(\log(\tfrac{n}{\gamma_{0}}))$ bits. Then from Weyl’s inequality $|\lambda_{k}(\widetilde{\boldsymbol{\mathrm{A}}})-\lambda_{k}(\boldsymbol{\mathrm{A}}\boldsymbol{\mathrm{A}}^{*}/4)|\leq\gamma_{0}/4$. We then call $\widetilde{\boldsymbol{\mathrm{X}}}\leftarrow\mathsf{REGULARIZE}(\widetilde{\boldsymbol{\mathrm{A}}},\gamma_{0},\delta_{0})$. Conditioning on success of Proposition 3.2, for all $i$ it holds that $|\lambda_{i}(\widetilde{\boldsymbol{\mathrm{X}}})-\lambda_{i}(\widetilde{\boldsymbol{\mathrm{A}}})|\leq 9\gamma_{0}/16$, and therefore $|\lambda_{i}(\widetilde{\boldsymbol{\mathrm{X}}})-\lambda_{i}(\boldsymbol{\mathrm{A}}\boldsymbol{\mathrm{A}}^{*}/4)|\leq\gamma_{0}$.
@@
Let $\lambda_{i}^{\prime}=\lambda_{i}(\widetilde{\boldsymbol{\mathrm{X}}})$ and $\lambda_{i}=\lambda_{i}(\boldsymbol{\mathrm{A}}\boldsymbol{\mathrm{A}}^{*}/4)$. Since all the $\lambda_{i}$ are non-negative, then all the $\lambda_{i}^{\prime}\in[-\gamma_{0},1+\gamma_{0}]$. We thus construct grid $\mathsf{g}=\operatorname{grid}(0,1+\gamma_{0},\gamma_{0})=\{0,1/8,2/8,3/8,\ldots,1,9/8\}$. We now need to perform counting queries similar to Theorem 3.1. Since we conditioned on success of Proposition 3.2, $\mathsf{REGULARIZE}$ ensures that for every $h_{j}\in\mathsf{g}$ it holds that $\sigma_{\min}(h_{j}-\widetilde{\boldsymbol{\mathrm{X}}}+\boldsymbol{\mathrm{E}}_{h_{j}})\geq\varepsilon_{0}$ for $\varepsilon_{0}=\frac{\gamma_{0}\delta_{0}}{8|\mathsf{g}|n\sqrt{\pi\ln(4n/\delta_{0})}}$. Then we execute $C_{j}\leftarrow\mathsf{COUNT}(\widetilde{\boldsymbol{\mathrm{X}}},h_{j},\varepsilon_{0})$ for every $h_{j}\in\mathsf{g}$ (in total $\Theta(1)$ calls to $\mathsf{COUNT}$). We set $j_{k}$ to be the smallest $j$ such that $C_{j}\geq k$. Now $\lambda^{\prime}_{1}\in[h_{j_{k}},h_{j_{k}+1}]$. Since $\lambda^{\prime}_{k}\in[\lambda_{k}-\gamma_{0},\lambda_{k}+\gamma_{0}]$, then $\lambda_{k}\in[h_{j_{1}-1}-\gamma_{0},h_{j_{k}}+\gamma_{0}]=[h_{j_{k}-2},h_{j_{k}+1}]:=I_{k}$, where $|I_{k}|=3\gamma_{0}$ (we can always ignore the negative portion of $I_{k}$, if there is any). In the next step we halve $\gamma_{0}$ to $\gamma_{1}=1/8$, and we create another grid inside $I_{k}$, $\mathsf{g}_{1}=\operatorname{grid}(h_{j_{k}-2},h_{j_{k}+1}+\gamma_{1},\gamma_{1})$. The new grid has at most $8$ points. We query all the points $h_{j}$ of $\mathsf{g}_{1}$ for $C_{j}\leftarrow\mathsf{COUNT}(\widetilde{\boldsymbol{\mathrm{X}}},h_{j},\epsilon_{1})$, and we pick a new $j_{1}=\arg\min\{h_{j}|C_{j}\geq k\}$. Now $\lambda_{k}$ lies inside a new interval $I_{k}$ with $|I_{k}|=3\gamma_{1}$, which is half the size of the interval of the previous iteration. If we set $\widetilde{\lambda}_{k}$ equal to the midpoint of $I_{k}$ (which is always positive), then $|\widetilde{\lambda}_{k}-\lambda_{k}|\leq 3\gamma/2$. We keep repeating the same procedure by halving $\gamma_{i}$ and $\delta_{i}$ in each step $i$.
@@
meaning that the algorithm will halt and return an $\epsilon$-accurate $\lambda_{k}$ in at most $O\left(\log(\tfrac{9}{\gamma_{m}})\right)=O\left(\log(\tfrac{1}{\epsilon\lambda_{k}})\right)$ iterations.
@@
arithmetic operations using $O\left(\log(n)\log(\tfrac{1}{\gamma_{i}\delta_{i}})\log^{3}(\tfrac{n}{\gamma_{i}\delta_{i}})\right)$ bits. The cost is maximized in the last iteration where $\gamma_{m}=\Theta(\epsilon\lambda_{k})$ and $\delta_{m}=\Theta(\delta_{0}\epsilon\lambda_{k})$, which gives
$O\left(T_{\mathsf{MM}}(n)\log(\tfrac{n}{\delta_{0}\epsilon\lambda_{k}})\log(\tfrac{1}{\epsilon\lambda_{k}})\right)$
@@
$O\left(\log(n)\log(\tfrac{1}{\delta_{0}\epsilon\lambda_{k}})\log^{3}(\tfrac{n}{\delta_{0}\epsilon\lambda_{k}})\right)$
@@
arithmetic operations and $O\left(\log(n)\log(\tfrac{1}{\delta\epsilon\sigma_{k}})\log^{3}(\tfrac{n}{\delta\epsilon\sigma_{k}})\right)$ bits. ∎
@@
Let $\boldsymbol{\mathrm{A}}\in\mathbb{H}^{n}_{++}$. Given $\epsilon\in(0,1)$ and $\delta\in(0,1/2)$, we can compute $\widetilde{\kappa}$ such that $\kappa(\boldsymbol{\mathrm{A}})\leq\widetilde{\kappa}\leq 32\kappa(\boldsymbol{\mathrm{A}})$, with an algorithm $\widetilde{\kappa}\leftarrow\mathsf{COND}(\boldsymbol{\mathrm{A}},\delta)$ in
@@
using $O\left(\log(n)\log^{3}(\tfrac{\kappa(\boldsymbol{\mathrm{A}})}{\delta})\log^{3}(\tfrac{n\kappa(\boldsymbol{\mathrm{A}})}{\delta})\right)$ bits of precision with probability at least $1-\delta$.

### v4

@@ B.1 Spectral gaps with diagonalization
The straightforward approach to compute the desired spectral gaps is to iteratively compute $\boldsymbol{\mathrm{S}}^{-1}\boldsymbol{\mathrm{H}}$ and diagonalize it, until the eigenvalues are well approximated. If we compute $\boldsymbol{\mathrm{A}}\leftarrow\mathsf{MM}(\mathsf{INV}(\boldsymbol{\mathrm{S}}),\boldsymbol{\mathrm{H}})$, then we can write $\boldsymbol{\mathrm{A}}=\boldsymbol{\mathrm{S}}^{-1}\boldsymbol{\mathrm{H}}+\boldsymbol{\mathrm{E}}$, for some error matrix $\boldsymbol{\mathrm{E}}$. The next step is to approximate $\operatorname{gap}_{k}(\boldsymbol{\mathrm{S}}^{-1}\boldsymbol{\mathrm{H}})$ and $\mu_{k}(\boldsymbol{\mathrm{S}}^{-1}\boldsymbol{\mathrm{H}})$. For this one could use the recent state-of-the-art backward-approximate diagonalization algorithm of [15].
Theorem B.1 ($\mathsf{EIG}$, imported Theorem 1.6 from [15]).
@@
Applying $\mathsf{EIG}$ on $\boldsymbol{\mathrm{A}}$ we can obtain a backward-approximate diagonalization. But we are not finished yet, since we are interested in each individual eigenvalue. To translate the backward error to a forward error for the eigenvalues, and, ultimately, the spectral gap, one can try to use Corollary 1.7 and Proposition 1.1 of [15]. However, this approach has two main limitations. First, it relies on simplicity of the spectrum, i.e., it assumes that the minimum gap between any pair of eigenvalues is larger than zero. This assumption is quite restrictive, since the desired gap might be well-defined even at the presence of other multiple eigenvalues. For example, in DFT applications it is not uncommon to have eigenenergies with algebraic multiplicity larger than one, and at the same time have a large band-gap that separates the occupied from the unoccupied orbitals. The second limitation is that the aforementioned Corollary 1.7 requires as an input parameter an actual over-estimate for the minimum eigenvalue gap. Even if such a gap exists, it is not described how to estimate it.
For diagonalizable matrices, we can leverage the following Corollary B.1, which is an immediate consequence of Kahan’s inequality (Fact A.1), and it overcomes the aforementioned limitations.
@@
Remark B.1.
We highlight that a similar result can be obtained by using the more recent pencil diagonalization algorithm of [41]. The authors the latter mention that their algorithm should be “more numerically stable” than $\mathsf{EIG}$ of [15], since it uses an inverse-free iteration internally, and they do provide strong theoretical and experimental evidence for this statement. However, as there is no formal, end-to-end proof for the bit complexity at the time of this writing, we choose to compare against $\mathsf{EIG}$ of [15]. Note that, any improvements on the bit requirements for the matrix sign function, directly provide the same improvements for our main Theorems 1.1 and 3.1.
@@ C.1 Error analysis
which means that it suffices to bound the individual terms in the sum. The first term is the error of the first recursive call. Following the notation of [39], let us denote $\boldsymbol{\mathsf{err}}(n)$ the norm-wise error of the algorithm for size $n$. We can then write
@@ C.2 Maintaining positive-definiteness, norms, and condition numbers throughout the recursion
C.3 Final backward-approximation bounds and proof of Theorem 1.2
Having safeguarded the possibility of a breakdown due to loss of positive-definiteness, we can now revisit the bounds of Lemma C.1 and finalize the proof of Theorem 1.2.
Theorem C.1 (Restatement of Theorem 1.2).
@@
The arithmetic complexity of the algorithm was already bounded in Proposition C.2.
For the error proof we assume the bound for $\textbf{{u}}\leq\textbf{{u}}_{++}$ from Inequality (13). Combining the inequalities for a single recursion step from Lemma C.2 with the results of the previous section and the bound for u, we have the following inequalities (recall once more that $\boldsymbol{\mathrm{M}}$ is the input of the algorithm for the specific recursive call, while $\boldsymbol{\mathrm{M}}_{0}$ is the matrix in the original call):
@@ C.4 Reducing the GEP to a regular Hermitian eigenproblem
Given $\boldsymbol{\mathrm{H}}\in\mathbb{H}^{n}$, $\boldsymbol{\mathrm{S}}\in\mathbb{H}^{n}_{++}$, and $\epsilon\in(0,1),$ Algorithm 4 executes $O(T_{\mathsf{MM}}(n))$ floating point operations and returns a matrix $\widetilde{\boldsymbol{\mathrm{H}}}\leftarrow\mathsf{REDUCE}(\boldsymbol{\mathrm{H}},\boldsymbol{\mathrm{S}},\epsilon)$, such that, if the machine precision satisfies
@@ C.5 Counterexample for floating point LU
In this section provide a counter-example which illustrates why the backward-stable LU factorization algorithm of [39] is not sufficient to obtain a Cholesky factor for a Hermitian postive-definite matrix. We apply one-by-one the steps of the $\mathsf{LUR}$ algorithm of [39] on the matrix:
@@ Appendix D Regularization with diagonal disorder
We now arrive to the more interesting part of the analysis. In Section 2 we argued that if we know the gap and the midpoint then the sign function can yield the desired spectral projector. To compute the gap and midpoint efficiently, we will take advantage of the symmetry induced by the Cholesky-based reduction, and use the Wegner estimate [138] for Hermitian diagonal perturbations to regularize the problem in the spirit of smoothed analysis. We use the following variant of the Wegner estimate.
@@
This directly comes from the well-known result of Wegner [138], which states that that for any interval $I\subset\mathbb{R}$
@@
Similar results can be obtained for other classes of random matrices $\boldsymbol{\mathrm{G}}$. Such an example is the Gaussian Unitary Ensemble (GUE) [4], in which case $\boldsymbol{\mathrm{G}}$ is dense, but it is invariant under rotations which might be useful for other applications.
@@ E.2 Approximate midpoint and gap
Theorem E.1 ($\mathsf{GAP}$, Restatement of Theorem 3.1).
@@
arithmetic operations using $O\left(\log(n)\left(\log^{4}(\tfrac{n}{\delta\epsilon\operatorname{gap}_{k}})+\log(\kappa(\boldsymbol{\mathrm{S}}))\right)\right)$ bits, where $\lambda_{i}$ are the eigenvalues of $(\boldsymbol{\mathrm{H}},\boldsymbol{\mathrm{S}})$. If $\kappa(\boldsymbol{\mathrm{S}})$ is unknown, additional $O(T_{\mathsf{MM}}(n)\log(\tfrac{n\kappa(\boldsymbol{\mathrm{S}})}{\delta})\log(\kappa(\boldsymbol{\mathrm{S}})))$ floating point operations and $O(\log(n)\log^{4}(\tfrac{n\kappa(\boldsymbol{\mathrm{S}})}{\delta}))$ bits are suficient to compute it with Corollary E.1.
@@
$O\left(T_{\mathsf{MM}}(n)\left(\log(\tfrac{1}{\varepsilon_{m}})+\log(\log(\tfrac{n}{\varepsilon_{m}}))\right)\right)=O\left(T_{\mathsf{MM}}(n)\log(\tfrac{n}{\gamma_{m}\delta_{m}})\right)$
arithmetic operations using $O\left(\log(n)\log^{3}(\tfrac{1}{\varepsilon_{m}})\log(\tfrac{n}{\varepsilon_{m}})\right)=O\left(\log(n)\log^{4}(\tfrac{n}{\gamma_{m}\delta_{m}})\right)$ bits. The last iteration where $\gamma_{m}=\Theta(\epsilon\operatorname{gap}_{k})$ and $\delta_{m}=\Theta(\delta_{0}\epsilon\operatorname{gap}_{k})$ is the most expensive, which gives a total of
$O\left(T_{\mathsf{MM}}(n)\log(\tfrac{n}{\delta_{0}\epsilon\operatorname{gap}_{k}})\cdot m\right)=O\left(T_{\mathsf{MM}}(n)\log(\tfrac{n}{\delta_{0}\epsilon\operatorname{gap}_{k}})\log(\tfrac{1}{\epsilon\operatorname{gap}_{k}})\right)$
@@
$O\left(\log(n)\log^{4}(\tfrac{n}{\delta_{0}\epsilon\operatorname{gap}_{k}})\right)$
@@
To be able to call $\mathsf{REDUCE}$ in every step, we need to approximate $\kappa(\boldsymbol{\mathrm{S}})$ in order to set the machine precision appropriately. For this we can use Corollary E.1, which is detailed in the next section, and it returns $\widetilde{\kappa}\in\Theta(\kappa(\boldsymbol{\mathrm{S}}))$ with probability $1-\delta$. It requires $O(T_{\mathsf{MM}}(n)\log(\tfrac{n\kappa(\boldsymbol{\mathrm{S}})}{\delta})\log(\kappa(\boldsymbol{\mathrm{S}})))$ floating point operations using $O(\log(n)\log^{4}(\tfrac{n\kappa(\boldsymbol{\mathrm{S}})}{\delta}))$ bits. ∎
@@ E.3 Singular values, singular gaps, and condition number
$O\left(T_{\mathsf{MM}}(n)\left(\tfrac{m}{n}+\log(\tfrac{n}{\delta\epsilon\sigma_{k}})\right)\log(\tfrac{1}{\epsilon\sigma_{k}})\right)$
floating point operations and requires $O\left(\log(n)\log^{4}(\tfrac{n}{\delta\epsilon\sigma_{k}})\right)$ bits.
@@
We will solve the problem with counting queries, similar to Theorem 3.1, halving $\gamma$ and $\delta$ in every iteration. We start with $\gamma_{0}=1/8$ and $\delta_{0}=\delta/2$. First, we divide $\boldsymbol{\mathrm{A}}$ by two, to ensure that $\|\boldsymbol{\mathrm{A}}\|\leq 1/2$. Then we construct $\widetilde{\boldsymbol{\mathrm{A}}}=\mathsf{HERM}(\mathsf{MM}(\boldsymbol{\mathrm{A}},\boldsymbol{\mathrm{A}}^{*}))=\boldsymbol{\mathrm{A}}\boldsymbol{\mathrm{A}}^{*}/4+\boldsymbol{\mathrm{E}}^{\mathsf{MM}}$. To perform this rectangular multiplication in practice, we partition $\boldsymbol{\mathrm{A}}$ in $m/n$ blocks of size $n\times n$ each, perform the individual multiplications and sum the results. We can ensure that $\|\boldsymbol{\mathrm{E}}^{\mathsf{MM}}\|\leq\gamma_{0}/4$ if we use $O(\log(\tfrac{n}{\gamma_{0}}))$ bits. Then from Weyl’s inequality $|\lambda_{k}(\widetilde{\boldsymbol{\mathrm{A}}})-\lambda_{k}(\boldsymbol{\mathrm{A}}\boldsymbol{\mathrm{A}}^{*}/4)|\leq\gamma_{0}/4$. We then call $\widetilde{\boldsymbol{\mathrm{X}}}\leftarrow\mathsf{REGULARIZE}(\widetilde{\boldsymbol{\mathrm{A}}},\gamma_{0},\delta_{0})$. Conditioning on success of Proposition 3.2, for all $i$ it holds that $|\lambda_{i}(\widetilde{\boldsymbol{\mathrm{X}}})-\lambda_{i}(\widetilde{\boldsymbol{\mathrm{A}}})|\leq 9\gamma_{0}/16$, and therefore $|\lambda_{i}(\widetilde{\boldsymbol{\mathrm{X}}})-\lambda_{i}(\boldsymbol{\mathrm{A}}\boldsymbol{\mathrm{A}}^{*}/4)|\leq\gamma_{0}$.
@@
Let $\lambda_{i}^{\prime}=\lambda_{i}(\widetilde{\boldsymbol{\mathrm{X}}})$ and $\lambda_{i}=\lambda_{i}(\boldsymbol{\mathrm{A}}\boldsymbol{\mathrm{A}}^{*}/4)$. Since all the $\lambda_{i}$ are non-negative, then all the $\lambda_{i}^{\prime}\in[-\gamma_{0},1+\gamma_{0}]$. We thus construct grid $\mathsf{g}=\operatorname{grid}(0,1+\gamma_{0},\gamma_{0})=\{0,1/8,2/8,3/8,\ldots,1,9/8\}$. We now need to perform counting queries similar to Theorem 3.1. Since we conditioned on success of Proposition 3.2, $\mathsf{REGULARIZE}$ ensures that for every $h_{j}\in\mathsf{g}$ it holds that $\sigma_{\min}(h_{j}-\widetilde{\boldsymbol{\mathrm{X}}}+\boldsymbol{\mathrm{E}}_{h_{j}})\geq\varepsilon_{0}$ for $\varepsilon_{0}=\frac{\gamma_{0}\delta_{0}}{8|\mathsf{g}|n\sqrt{\pi\ln(4n/\delta_{0})}}$. Then we execute $C_{j}\leftarrow\mathsf{COUNT}(\widetilde{\boldsymbol{\mathrm{X}}},h_{j},\varepsilon_{0})$ for every $h_{j}\in\mathsf{g}$ (in total $\Theta(1)$ calls to $\mathsf{COUNT}$). We set $j_{k}$ to be the smallest $j$ such that $C_{j}\geq k$. Now $\lambda^{\prime}_{1}\in[h_{j_{k}},h_{j_{k}+1}]$. Since $\lambda^{\prime}_{k}\in[\lambda_{k}-\gamma_{0},\lambda_{k}+\gamma_{0}]$, then $\lambda_{k}\in[h_{j_{1}-1}-\gamma_{0},h_{j_{k}}+\gamma_{0}]=[h_{j_{k}-2},h_{j_{k}+1}]:=I_{k}$, where $|I_{k}|=3\gamma_{0}$ (we can always ignore the negative portion of $I_{k}$, if there is any).
In the next step we halve $\gamma_{0}$ to $\gamma_{1}=1/8$, and we create another grid inside $I_{k}$, $\mathsf{g}_{1}=\operatorname{grid}(h_{j_{k}-2},h_{j_{k}+1}+\gamma_{1},\gamma_{1})$. The new grid has at most $8$ points. We query all the points $h_{j}$ of $\mathsf{g}_{1}$ for $C_{j}\leftarrow\mathsf{COUNT}(\widetilde{\boldsymbol{\mathrm{X}}},h_{j},\epsilon_{1})$, and we pick a new $j_{1}=\arg\min\{h_{j}|C_{j}\geq k\}$. Now $\lambda_{k}$ lies inside a new interval $I_{k}$ with $|I_{k}|=3\gamma_{1}$, which is half the size of the interval of the previous iteration. If we set $\widetilde{\lambda}_{k}$ equal to the midpoint of $I_{k}$ (which is always positive), then $|\widetilde{\lambda}_{k}-\lambda_{k}|\leq 3\gamma/2$. We keep repeating the same procedure by halving $\gamma_{i}$ and $\delta_{i}$ in each step $i$.
@@
meaning that the algorithm will halt and return an $\epsilon$-accurate $\lambda_{k}$ in at most $m=O\left(\log(\tfrac{9}{\gamma_{m}})\right)=O\left(\log(\tfrac{1}{\epsilon\lambda_{k}})\right)$ iterations.
@@
arithmetic operations using $O\left(\log(n)\log^{4}(\tfrac{n}{\gamma_{i}\delta_{i}})\right)$ bits. The cost is maximized in the last iteration where $\gamma_{m}=\Theta(\epsilon\lambda_{k})$ and $\delta_{m}=\Theta(\delta_{0}\epsilon\lambda_{k})$, which gives
$O\left(T_{\mathsf{MM}}(n)\log(\tfrac{n}{\delta_{0}\epsilon\lambda_{k}})\cdot m\right)=O\left(T_{\mathsf{MM}}(n)\log(\tfrac{n}{\delta_{0}\epsilon\lambda_{k}})\log(\tfrac{1}{\epsilon\lambda_{k}})\right)$
@@
$O\left(\log(n)\log^{4}(\tfrac{n}{\delta_{0}\epsilon\lambda_{k}})\right)$
@@
arithmetic operations and $O\left(\log(n)\log^{4}(\tfrac{n}{\delta\epsilon\sigma_{k}})\right)$ bits. ∎
@@
Let $\boldsymbol{\mathrm{A}}\in\mathbb{H}^{n}_{++}$. Given $\delta\in(0,1/2)$, we can compute $\widetilde{\kappa}$ such that $\kappa(\boldsymbol{\mathrm{A}})\leq\widetilde{\kappa}\leq 32\kappa(\boldsymbol{\mathrm{A}})$, with an algorithm $\widetilde{\kappa}\leftarrow\mathsf{COND}(\boldsymbol{\mathrm{A}},\delta)$ in
@@
using $O\left(\log(n)\log^{4}(\tfrac{n\kappa(\boldsymbol{\mathrm{A}})}{\delta})\right)$ bits of precision with probability at least $1-\delta$.